OpenStudy (anonymous):

Sara's mom takes the parkway to work. It takes her 30 minutes to travel from mile marker 88 where she enters to her exit at mile marker 121. What is her average speed? Solve?

Directrix (directrix):

121- 88 = 33 miles traversed in 30 minutes. d = r*t 33 = r* (.5 hr) Solve for r. Average speed will be in miles per hour. @6unicorns

OpenStudy (anonymous):

thanks

Directrix (directrix):

You're welcome.